OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (hoblos):

4x - 3y = -12 3y= 4x +12 y = 4/3 x +4 so the slope is 4/3 y-4=4/3(x+3) y= 4/3x +8